### v1

@@ Abstract
Pneumonia, a severe respiratory disease, poses significant diagnostic challenges, especially in underdeveloped regions. Traditional diagnostic methods, such as chest X-rays, suffer from variability in interpretation among radiologists, necessitating reliable automated tools. In this study, we propose a novel approach combining deep learning and transformer-based attention mechanisms to enhance pneumonia detection from chest X-rays. Our method begins with lung segmentation using a TransUNet model that integrates our specialized transformer module, which has fewer parameters compared to common transformers while maintaining performance. This model is trained on the ”Chest Xray Masks and Labels” dataset and then applied to the Kermany and Cohen datasets to isolate lung regions, enhancing subsequent classification tasks. For classification, we employ pre-trained ResNet models (ResNet-50 and ResNet-101) to extract multi-scale feature maps, processed through our modified transformer module. By employing our specialized transformer, we attain superior results with significantly fewer parameters compared to common transformer models. Our approach achieves high accuracy rates of 92.79% on the Kermany dataset and 95.11% on the Cohen dataset, ensuring robust and efficient performance suitable for resource-constrained environments. https://github.com/amirrezafateh/Multi-Scale-Transformer-Pneumonia
@@ 1 Introduction
In order to effectively treat pneumonia and improve patient outcomes, early diagnosis is essential. Still, identifying pneumonia can be challenging. It can be easily confused with other lung diseases, and interpreting X-ray images can vary significantly among radiologists, making consistent and accurate diagnosis difficult health_poverty_action_poverty_and_health . Conventional imaging methods are frequently employed, including computed tomography (CT), magnetic resonance imaging (MRI), and chest radiographs. Chest X-rays are favored for being non-invasive and cost-effective radiologyinfo_chest_xray . However, the variability in interpreting these images makes more reliable, automated diagnostic tools necessary.
The advent of deep learning technology presents a promising approach to overcoming the difficulties in diagnosing pneumonia gayathri2022computer ; lin2023enhancing ; ortiz2022automatic . As a branch of artificial intelligence (AI), deep learning entails training neural networks on extensive datasets to identify patterns and generate predictions. In medical imaging, deep learning algorithms can scrutinize chest X-rays with remarkable precision, often exceeding human radiologists in terms of consistency and speed. Convolutional Neural Networks (CNNs), a specialized deep learning model designed for image analysis, have been developed to identify pneumonia from chest X-rays. These models can identify subtle features indicative of pneumonia that might be missed by the human eye, achieving diagnostic performance on par with, and sometimes exceeding, that of experienced radiologists krizhevsky2012imagenet ; parhami2023comparison . Recent advancements include the use of ensemble learning techniques, which combine predictions from multiple models to improve overall accuracy, and the integration of attention mechanisms, which help models focus on the most relevant parts of the image bhatt2023convolutional ; gu2020net ; fateh2024advancing . Moreover, initially designed for natural language processing, transformers have demonstrated significant potential in visual tasks. They excel at modeling long-range dependencies and identifying intricate patterns in medical images park2022multi ; mabrouk2022pneumonia .
@@
A significant challenge in training these deep learning models is the requirement for large, high-quality labeled datasets, which are often hard to acquire in the medical field. Transfer learning tackles this problem by utilizing models pre-trained on extensive datasets like ImageNet and then fine-tuning them for particular medical tasks using smaller datasets. This approach has significantly improved the accuracy of medical image classifications chen2017rethinking . Despite the benefits of using transformers and their achievements in segmentation and classification tasks in CXR images, they bring two new challenges: models are more complex and have many parameters. These complex models used more hardware resources and took longer to train.
We propose an innovative method combining deep learning with transformer-based attention mechanisms to overcome these challenges while maintaining lower model complexity. Our method begins with lung segmentation using a TransUNet model, which integrates transformer-based attention mechanisms into the U-Net architecture. The TransUNet model is trained on the ”Chest Xray Masks and Labels” dataset candemir2013lung ; jaeger2013automatic to accurately segment lung regions in the images. Once trained, this pre-trained model is used with frozen weights to predict lung masks for our target datasets, Kermany kermany2018labeled and Cohen cohen2020covid . This segmentation step isolates the lung regions, enhancing the subsequent classification task.
For classification, we utilize pre-trained ResNet models, specifically ResNet-50 and ResNet-101, as the foundation for feature extraction. We draw multi-scale feature maps from various stages of the ResNet models and process them through a customized transformer module that uses a cross-attention mechanism. This transformer has been optimized to minimize the number of parameters while preserving performance. By concentrating on the relevant lung regions and integrating multi-scale information, our approach strives to achieve high diagnostic accuracy for pneumonia detection. This architecture reduces the computational load and ensures robust and reliable performance, making it suitable for deployment in resource-limited settings.
@@
  • Achieving high accuracy rates of 92.79% on the Kermany dataset and 95.11% on the Cohen dataset.
@@ 2 Related Work
In recent years, the focus of research on diagnosing and categorizing lung diseases, including pneumonia, through medical imaging has intensified, driven by advances in machine learning and deep learning technologies. Precisely segmenting lung areas in chest X-ray (CXR) images is essential for reliable disease identification and thorough analysis. This section explores deep learning techniques for segmenting and diagnosing lung diseases in chest X-ray (CXR) images. It focuses on the U-Net architecture and its variations, including attention mechanisms and transformer blocks, which have revolutionized lung disease segmentation. The approaches are categorized into basic Deep Learning models, transfer learning, fine-tuning, and custom models, highlighting the progression and impact of these advanced techniques on improving diagnostic outcomes.
@@ 2.2.2 Deep Learning Models
The beginning of deep learning, especially Convolutional Neural Networks (CNNs), has significantly transformed medical image analysis by providing superior accuracy and robustness. For instance, Stephen et al. designed a custom CNN model from scratch, achieving a training accuracy of 95.31% and a validation accuracy of 93.73% stephen2019efficient . Similarly, Sharma et al. created a straightforward CNN architecture that reached a 90.68% accuracy rate on the Kermany dataset using data augmentation sharma2020feature ; kermany2018labeled . However, relying solely on data augmentation does not introduce substantially new information, restricting the model’s ability to learn a wide range of complex patterns from the training data.
@@ 3.2.3 Applying the Trained TransUNet to the Cohen Dataset
After successfully training the TransUNet model on the ”Chest Xray Masks and Labels” dataset, we utilize this pre-trained segmentation model to predict lung masks for the Cohen dataset. This step is crucial for enhancing the accuracy of the subsequent classification task by focusing on the lung regions within the X-ray images.
The Cohen dataset, which contains chest X-ray images, requires preprocessing to ensure that our classification model focuses on the most relevant regions. To achieve this, we apply the trained TransUNet model to segment the lung areas from these images. The Cohen dataset is first preprocessed to match the input requirements of the TransUNet model. This involves standardizing the image dimensions and normalizing the pixel values to ensure consistency with the training data used for the TransUNet model.
Using the pre-trained TransUNet model, we generate lung masks for each X-ray image in the Cohen dataset. The segmentation model outputs binary masks that highlight the lung regions while suppressing the background. By utilizing the pre-trained TransUNet model to segment the lung regions in the Cohen dataset, we effectively preprocess the data to improve the performance of our classification model. This segmentation step filters out noise and irrelevant features, allowing the classifier to concentrate on the lung areas, thereby enhancing the overall accuracy and robustness of our proposed method.
@@ 3.3.1 Backbone
The backbone of our proposed method utilizes a pre-trained ResNet model, specifically ResNet-50 or ResNet-101, to extract multi-scale feature maps from the input chest X-ray images. Initially, we resize all images to 512x512 pixels. We focus on the outputs from Blocks 2, 3, and 4 of the ResNet, denoted as $B^{2}$, $B^{3}$, and $B^{4}$ respectively.ach of these blocks provides feature maps of size $(c,h,w)$, where $h=w=64$ The channels $c$ for these blocks are 512, 1024, and 2048 respectively.
@@ 3.3.2 Transformer
The feature map $B^{\prime 4}$ is reshaped to form the key and value inputs for the transformer. Specifically, as shown in equation 6, $B^{\prime 4}$ is reshaped from (c,h,w) to (c,hxw).
@@ 4.1 Dataset
For classification task, we used two datasets. COVID-19 Image Data Collection provided by Cohen cohen2020covid . This dataset comprises a total of 6,432 images, including three classes: Pneumonia, COVID-19, and Normal. The dataset is notably challenging due to its class imbalance and the complexity introduced by the three distinct classes. The distribution of images in the training set is as follows: 3,418 images of Pneumonia, 1,266 images of Normal, and 460 images of COVID-19. Approximately 20% of the images are allocated for testing.
@@ 4.2 Experimental Setting
Table 1: Performance on cohen dataset. Numbers in bold represent the best performance, while underlined values denote the second-best performance.
@@
Densenet121 87.8% 53.9% 71.0% 61.27%
Densenet169 87.1% 32.3% 65.6% 43.28%
Densenet201 88.4% 51.9% 79.0% 62.64%
Mobilenet_v2 86.9% 33.4% 75.0% 46.21%
ResNet-50 87.1% 38.4% 71.0% 49.84%
ResNet-101 87.9% 33.5% 73.0% 45.92 %
@@
Proposed method (ResNet-50 as backbone) 95.11% 95.81% 94.04% 94.89%
Proposed method (ResNet-101 as backbone) 94.18% 94.76% 92.79% 93.72%
Table 2: Performance on kermany dataset. Numbers in bold represent the best performance, while underlined values denote the second-best performance.
@@
Table 3: effect of segmentation on cohen dataset
@@
ResNet-50 on original images 91.23% 90.94% 85.60% 87.98% 89.41%
on predicted masks 95.11% 95.81% 94.04% 94.89% 95.07%
ReNet101 on original images 90.22% 88.16% 87.04% 87.54% 90.16%
on predicted masks 94.18% 94.76% 92.79% 93.72% 94.28%
@@ 4.3 Evaluation Metrics
Figure 3: confusion matrix of cohen dataset with different scenarios
@@ 4.5 Ablation study
To evaluate the impact of the segmentation component on the performance of our classification model, we conducted an ablation study using the Cohen dataset. This study compares the classification results obtained with and without the segmentation step, providing insights into the effectiveness of incorporating lung masks generated by the TransUNet model.
The ablation study involves evaluating the classification performance of our model with two different input scenarios: 1) Original Images: Classification is performed directly on the raw chest X-ray images from the Cohen dataset. 2) Predicted Masks: Classification is performed on the chest X-ray images after segmenting the lung regions using the TransUNet model. The images used for classification are limited to the areas highlighted by the predicted lung masks.
@@ 5 Conclusion
This paper presents an innovative and efficient method for pneumonia detection utilizing a novel multi-scale transformer approach. By integrating lung segmentation using the TransUNet model with a specialized transformer module, our approach effectively isolates lung regions, thereby enhancing the performance of subsequent classification tasks. The proposed method demonstrates significant improvements in classification metrics, as evidenced by the ablation study on the Cohen dataset. Both ResNet-50 and ResNet-101 backbones benefited from the segmentation masks, showing increased accuracy, precision, recall, and F1-score. These improvements underscore the effectiveness of our approach in focusing on relevant lung features while reducing noise from irrelevant regions.
The high accuracy rates of 92.79% on the Kermany dataset and 95.11% on the Cohen dataset confirm the robustness and reliability of our model. The reduction in the number of parameters compared to other state-of-the-art transformer models highlights our contribution to creating a more efficient yet powerful diagnostic tool suitable for deployment in resource-constrained environments. Our work paves the way for future research in several areas. Future work could explore further optimization of the transformer module to enhance performance and reduce computational complexity. Additionally, expanding the dataset to include a broader variety of pneumonia cases and other respiratory diseases could improve the model’s generalization.

### v2

@@ Abstract
Pneumonia, a widespread respiratory infection, remains a leading cause of morbidity and mortality globally, especially in vulnerable populations. Chest X-rays are a primary tool in detecting pneumonia; however, differences in imaging conditions and subtle visual indicators can make consistent interpretation challenging. Automated tools can complement traditional methods by providing additional insights to enhance diagnostic reliability and support clinical decision-making. In this study, we propose a novel approach that leverages deep learning with integrated self-attention mechanisms to enhance pneumonia detection in chest X-rays. Our method begins with lung segmentation using a TransUNet model that integrates our specialized transformer module, which has fewer parameters compared to common transformers while maintaining performance. This model is trained on the ”Chest Xray Masks and Labels” dataset and then applied to the ”Kermany” and ”Cohen” datasets to generate lung regions, enhancing subsequent classification tasks. By isolating the lung areas, the segmentation step provides focused input for the classification stage, enabling the model to concentrate on relevant regions. For classification, we utilize pre-trained ResNet models (ResNet-50 and ResNet-101) specifically to extract multi-scale feature maps, a core aspect of our approach. These feature maps are then processed through our modified transformer module to enhance pneumonia detection. By employing our specialized transformer, we achieve superior results with significantly fewer parameters compared to common transformer models. Our approach achieves high accuracy rates of 92.79% on the ”Kermany” dataset and 95.11% on the ”Cohen” dataset, ensuring robust and efficient performance suitable for resource-constrained environments. https://github.com/amirrezafateh/Multi-Scale-Transformer-Pneumonia
@@ 1 Introduction
In order to effectively treat pneumonia and improve patient outcomes, early diagnosis is essential. Still, identifying pneumonia can be challenging. It can be easily confused with other lung diseases, and making consistent and accurate diagnosis difficult health_poverty_action_poverty_and_health . Conventional imaging methods are frequently employed, including computed tomography (CT), magnetic resonance imaging (MRI), and chest radiographs. But, chest X-rays are favored for being non-invasive and cost-effective radiologyinfo_chest_xray .
The variability in interpreting chest X-ray images underscores the need for consistent, automated diagnostic tools. The development of deep learning technology offers a promising approach to addressing the challenges in pneumonia diagnosis gayathri2022computer ; lin2023enhancing ; ortiz2022automatic . As a branch of artificial intelligence (AI), deep learning entails training neural networks on extensive datasets to identify patterns and generate predictions. In medical imaging, deep learning algorithms can scrutinize chest X-rays with remarkable precision, often exceeding human radiologists in terms of consistency and speed. Convolutional Neural Networks (CNNs), a specialized deep learning model designed for image analysis, have been developed to identify pneumonia from chest X-rays askari2024enhancing . These models are trained to recognize patterns and subtle features associated with pneumonia, contributing to consistent and reliable diagnostic performance that complements the expertise of radiologists krizhevsky2012imagenet ; parhami2023comparison . However, traditional CNN-based models still cannot achieve impressive accuracy. Recent advancements include the use of ensemble learning techniques, which combine predictions from multiple models to improve overall accuracy, and the integration of attention mechanisms, which help models focus on the most relevant parts of the image fateh2024advancing . Transformers have demonstrated significant potential in visual tasks. They excel at modeling long-range dependencies and identifying intricate patterns in medical images fateh2024msdnet ; mabrouk2022pneumonia .
@@
A significant challenge in training these deep learning models is the requirement for large, high-quality labeled datasets, which are often hard to acquire in the medical field. Transfer learning tackles this problem by utilizing models pre-trained on extensive datasets like ImageNet and then fine-tuning them for particular medical tasks using smaller datasets. This approach has significantly improved the accuracy of medical image classifications chen2017rethinking . Even with transfer learning, models trained in one domain often fail to perform well in significantly different domains, such as medical applications. This highlights the need for more advanced and higher-level features, which is why transformers are increasingly used rezvani2024fusionlungnet . Despite the benefits of using transformers and their achievements in segmentation and classification tasks in CXR images, they bring two new challenges: models are more complex and have many parameters. These complex models used more hardware resources and took longer to train.
We propose an innovation approach that leverages deep learning with integrated self-attention mechanisms to overcome the challenges of traditional transformers while maintaining lower model complexity. Our method begins with lung segmentation using a TransUNet model, which integrates transformer-based attention mechanisms into the U-Net architecture. The TransUNet model is trained on the ”Chest Xray Masks and Labels” dataset candemir2013lung ; jaeger2013automatic to accurately segment lung regions in the images. Once trained, this pre-trained model is used with frozen weights to predict lung masks for our target datasets, ”Kermany” kermany2018labeled and ”Cohen” cohen2020covid . This segmentation step isolates the lung regions, enhancing the subsequent classification task.
For classification, we utilize pre-trained ResNet models, specifically ResNet-50 and ResNet-101, as the foundation for feature extraction. By drawing multi-scale feature maps from various stages of the ResNet models, we can leverage multiple feature spaces, which enhances the accuracy of our detection. This is achieved through a customized transformer module that employs a cross-attention mechanism, allowing us to make decisions based on more than one feature space. This transformer has been optimized to minimize the number of parameters while preserving performance. By concentrating on the relevant lung regions and integrating multi-scale information, our approach strives to achieve high diagnostic accuracy for pneumonia detection. This architecture reduces the computational load and ensures robust and reliable performance, making it suitable for deployment in resource-limited settings.
@@
  • Incorporation of multi-scale feature extraction, enabling enhanced performance through the utilization of multiple feature spaces.
@@
  • Achieving high accuracy rates of 92.79% on the ”Kermany” dataset and 95.11% on the ”Cohen” dataset.
@@ 2 Related Work
In recent years, the focus of research on diagnosing and categorizing lung diseases, including pneumonia, through medical imaging has intensified, driven by advances in machine learning and deep learning technologies. Precisely segmenting lung areas in chest X-ray (CXR) images is essential for reliable disease identification and thorough analysis. This section examines deep learning techniques for segmenting and diagnosing lung diseases in chest X-ray (CXR) images. For the segmentation task, we focus on the U-Net architecture and its variations, including attention mechanisms and transformer blocks, which have significantly advanced lung disease segmentation. In the classification task, we categorize approaches into basic deep learning models, transfer learning, fine-tuning, and custom models, emphasizing how these advanced techniques have progressively improved diagnostic outcomes.
@@ 2.2.2 Deep Learning Models
The beginning of deep learning, especially Convolutional Neural Networks (CNNs), has significantly transformed medical image analysis by providing superior accuracy and robustness. For instance, Stephen et al. designed a custom CNN model from scratch, achieving a training accuracy of 95.31% and a validation accuracy of 93.73% stephen2019efficient . Similarly, Sharma et al. created a straightforward CNN architecture that reached a 90.68% accuracy rate on the ”Kermany” dataset using data augmentation sharma2020feature ; kermany2018labeled . However, relying solely on data augmentation does not introduce substantially new information, restricting the model’s ability to learn a wide range of complex patterns from the training data.
@@ 3.2.3 Applying the Trained TransUNet to the Cohen Dataset
After successfully training the TransUNet model on the ”Chest Xray Masks and Labels” dataset, we utilize this pre-trained segmentation model to predict lung masks for the ”Cohen” dataset. This step is crucial for enhancing the accuracy of the subsequent classification task by focusing on the lung regions within the X-ray images.
The ”Cohen” dataset, which contains chest X-ray images, requires preprocessing to ensure that our classification model focuses on the most relevant regions. To achieve this, we apply the trained TransUNet model to segment the lung areas from these images. The ”Cohen” dataset is first preprocessed to match the input requirements of the TransUNet model. This involves standardizing the image dimensions and normalizing the pixel values to ensure consistency with the training data used for the TransUNet model.
Using the pre-trained TransUNet model, we generate lung masks for each X-ray image in the ”Cohen” dataset. The segmentation model outputs binary masks that highlight the lung regions while suppressing the background. By utilizing the pre-trained TransUNet model to segment the lung regions in the ”Cohen” dataset, we effectively preprocess the data to improve the performance of our classification model. This segmentation step filters out noise and irrelevant features, allowing the classifier to concentrate on the lung areas, thereby enhancing the overall accuracy and robustness of our proposed method.
@@ 3.3.1 Backbone
The backbone of our proposed method utilizes a pre-trained ResNet model, specifically ResNet-50 or ResNet-101, to extract multi-scale feature maps from the input chest X-ray images. Initially, we resize all images to 512x512 pixels. We focus on the outputs from Blocks 2, 3, and 4 of the ResNet, denoted as $B^{2}$, $B^{3}$, and $B^{4}$ respectively. Each of these blocks provides feature maps of size $(c,h,w)$, where $h=w=64$ The channels $c$ for these blocks are 512, 1024, and 2048 respectively.
@@ 3.3.2 Transformer
The feature map $B^{\prime 4}$ is reshaped to form the key and value inputs for the transformer. Specifically, as shown in equation 6, $B^{\prime 4}$ is reshaped from $(c,h,w)$ to $(c,h\times w)$.
@@ 4.1 Dataset
For classification task, we used two datasets. COVID-19 Image Data Collection provided by ”Cohen” cohen2020covid . This dataset comprises a total of 6,432 images, including three classes: Pneumonia, COVID-19, and Normal. The dataset is notably challenging due to its class imbalance and the complexity introduced by the three distinct classes. The distribution of images in the training set is as follows: 3,418 images of Pneumonia, 1,266 images of Normal, and 460 images of COVID-19. Approximately 20% of the images are allocated for testing.
@@ 4.2 Experimental Setting
Table 1: Performance on ”Cohen” dataset. Numbers in bold represent the best performance, while underlined values denote the second-best performance.
@@
Densenet121* 87.8% 53.9% 71.0% 61.27%
Densenet169* 87.1% 32.3% 65.6% 43.28%
Densenet201* 88.4% 51.9% 79.0% 62.64%
Mobilenet_v2* 86.9% 33.4% 75.0% 46.21%
ResNet-50* 87.1% 38.4% 71.0% 49.84%
ResNet-101* 87.9% 33.5% 73.0% 45.92 %
@@
Proposed method (ResNet-50 as backbone) 95.11% 95.81% 94.04% 94.92%
Proposed method (ResNet-101 as backbone) 94.18% 94.76% 92.79% 93.76%
Models marked with ”*” have results directly reported from goodwin2020intra .
Table 2: Performance on ”Kermany” dataset. Numbers in bold represent the best performance, while underlined values denote the second-best performance.
@@
Table 3: Effect of segmentation on ”Cohen” dataset
@@
ResNet-50 on original images 91.23% 90.94% 85.60% 88.19% 89.41%
on predicted masks 95.11% 95.81% 94.04% 94.92% 95.07%
ReNet101 on original images 90.22% 88.16% 87.04% 87.6% 90.16%
on predicted masks 94.18% 94.76% 92.79% 93.76% 94.28%
@@ 4.3 Evaluation Metrics
Figure 3: Confusion matrix of ”Cohen” dataset with different scenarios
@@ 4.5 Ablation study
To evaluate the impact of the segmentation component on the performance of our classification model, we conducted an ablation study using the ”Cohen” dataset. This study compares the classification results obtained with and without the segmentation step, providing insights into the effectiveness of incorporating lung masks generated by the TransUNet model.
The ablation study involves evaluating the classification performance of our model with two different input scenarios: 1) Original Images: Classification is performed directly on the raw chest X-ray images from the ”Cohen” dataset. 2) Predicted Masks: Classification is performed on the chest X-ray images after segmenting the lung regions using the TransUNet model. The images used for classification are limited to the areas highlighted by the predicted lung masks.
@@
Table 4: The effect of using each key component on ”Cohen” dataset (on original images and without segmentation)
Backbones Baseline
Multi-scale
feature maps
Transformer Accuracy Precision Recall F1-score
ResNet-50 X 84.62% 75.68% 75.10% 75.39%
X X 87.73% 80.55% 79.62% 80.08%
X X 87.73% 80.63% 80.47% 80.55%
X X X 91.23% 90.94% 85.60% 88.19%
ResNet-101 X 83.93% 74.21% 73.92% 74.06%
X X 85.56% 78.21% 80.4% 79.29%
X X 85.71% 78.59% 81.11% 79.83%
X X X 90.22% 88.16% 87.04% 87.6%
@@
Furthermore, we investigate the contributions of key components within our model on the ”Cohen” dataset, specifically focusing on the impact of multi-scale feature maps and the transformer module. The results are presented in Table 4, highlighting the performance metrics, including accuracy, precision, recall, and F1-score, across different configurations of the ResNet-50 and ResNet-101 backbones. Each row of the table demonstrates how the incorporation of each component affects the model’s performance, providing a comprehensive view of their individual and combined effects.
The findings reveal a significant enhancement in model performance when both multi-scale feature maps and the transformer are employed alongside the baseline configuration. For instance, with the ResNet-50 backbone, the accuracy improves from 84.62% (baseline) to 91.23% when all components are utilized. Similarly, the ResNet-101 backbone exhibits a notable increase in accuracy from 83.93% to 90.22%. These results underscore the effectiveness of our proposed innovations, illustrating that the integration of multi-scale feature maps and transformer elements not only enhances overall accuracy but also boosts precision, recall, and F1-score, which are crucial for the reliability of classification tasks. This highlights the importance of these key components in achieving improved performance in deep learning models for image analysis.
@@ 5 Conclusion
This paper presents an innovative and efficient method for pneumonia detection utilizing a novel multi-scale transformer approach. By integrating lung segmentation using the TransUNet model with a specialized transformer module, our approach effectively isolates lung regions, thereby enhancing the performance of subsequent classification tasks. The proposed method demonstrates significant improvements in classification metrics, as evidenced by the ablation study on the ”Cohen” dataset. Both ResNet-50 and ResNet-101 backbones benefited from the segmentation masks, showing increased accuracy, precision, recall, and F1-score. These improvements underscore the effectiveness of our approach in focusing on relevant lung features while reducing noise from irrelevant regions.
The high accuracy rates of 92.79% on the ”Kermany” dataset and 95.11% on the ”Cohen” dataset confirm the robustness and reliability of our model. The reduction in the number of parameters compared to other state-of-the-art transformer models highlights our contribution to creating a more efficient yet powerful diagnostic tool suitable for deployment in resource-constrained environments. Our work paves the way for future research in several areas. Future work could explore further optimization of the transformer module to enhance performance and reduce computational complexity. Additionally, expanding the dataset to include a broader variety of pneumonia cases and other respiratory diseases could improve the model’s generalization.